Change each logarithmic form to an equivalent exponential form:
(please show your work on how you reached the deteriminations)
a. log( 3, 81) = 4
b. log( 10, 0.0001) = 4
c. log( 3, (1/243) = -5
--------------------------------------------
I'm not sure if the formula above turned out correctly, but if it didn't, I tried to spell it out below:

a. Should read, log with the 3 below the g, then 81 =4

b. Should read, log with the 10 below the g, then 0.0001 = –4

c. Should read log with the 3 below the g, then in paranthesis 1 over 243,
end of paranthesis = - 5


Answer by stanbon(48535)

Change each logarithmic form to an equivalent exponential form:
(please show your work on how you reached the deteriminations)
a. log( 3, 81) = 4
b. log( 10, 0.0001) = -4
c. log( 3, (1/243) = -5
--------------------------------------------
Comment: log means "exponent of"
a. Your problem says "the exponent of 3 is 4 ; the result is 81"
3^4 = 81
----------------------------------
b. Your problem says "the exponent of 10 is -4 ;the result is 0.0001"
10^-4 = 0.0001
----------------------------------
c. Your problem says "the exponent of 3 is -5 ; the result is (1/243)"
3^-5 = 1/243
----------------------------------------
Cheers,
Stan H.